STUDY OF THE A^{1}Σ_{u}^{+} AND b^{3}Π_{0u} STATES IN Cs_{2} : NEW DATA AND GLOBAL ANALYSIS

\underline{HOUSSAM SALAMI}, THOMAS BERGEMAN, Department of Physics and Astronomy, SUNY, Stony Brook, NY 11794-3800; OLIVIER DULIEU, Laboratoire Aimé Cotton, CNRS, Bât. 505, Campus d’Orsay, 91405 Orsay France; DAN LI, FENG XIE and LI LI, Department of Physics, Tsinghua University, Beijing 100084, China.

The lowest excited states of alkali dimers are of particular interest in the creation of ultracold molecules and as pathways to higher states. Levels with mixed singlet-triplet character receive also considerable attention as window levels. The study of the A^{1}Σ_{u}^{+} and b^{3}Π_{0u^{+}} states in the heavier alkali molecules is relatively complicated due to the large spin-orbit interactions which mix the potentials of these two states. New, low-resolution data on the b^{3}Π_{0u^{±}} states of Cs_{2} have been obtained recently in Tsinghua University. These data characterize the bottom of the potential well of the b state, provide vibrational assignment and determine the value of Te, and allow us to analyze and fit previously obtained data form Lab. Aimé Cotton on mixed A/b states^{a}. The complete set of observations has been modeled using the discrete variable representation. Potential energy curves have been calculated relying on ab initio potentials and Morse-type spin-orbit functions. Currently, the rms residual of our fit to the new data is 1.65 cm^{−1} (∼ the experimental uncertainty of the dispersed fluorescence recorded on a monochromator) and 0.14 cm^{−1} to the older Fourier transform spectroscopy data which has experimental uncertainties of 0.003 cm^{−1}. Additional data would be very valuable, both in the regions where no levels have yet been observed and in those where few levels with dominant triplet character have been identified.

Work at Stony Brook was supported by NSF grant PHY 0652459.

aJ. Vergès and C. Amiot, J. Mol. Spectrosc. \underline{126} 393 (1987), C. Amiot and O. Dulieu, J. Chem. Phys. \underline{117} 5155 (2002)